OpenStudy (anonymous):

Find the radian measure of an angle of -280degrees

OpenStudy (anonymous):

a. 9/-14pi b. -14pi/9 *My answer c. 9pi/-14 d. -14/9pi

OpenStudy (anonymous):

2 pi = 360 degrees. Does that help?

OpenStudy (anonymous):

Um not really... sorry

OpenStudy (anonymous):

This is the conversion factors it shows how many radians are equal to one degree.

OpenStudy (sweetburger):

well pi would never be in the denominator so eliminate 2 of them...

OpenStudy (anonymous):

You have -280 degrees. So to get -280 to radians you multiplly by -280(2pi/360)

OpenStudy (anonymous):

Soo.. B?

OpenStudy (anonymous):

The reason why I did that is because 280 is in degrees and 360 is also in degrees. So, this would also cancel out as they are divide by each other.

OpenStudy (anonymous):

Im sorry but I am so so confused

OpenStudy (wolf1728):

There are 2 PI radians per 360° So in -280° there are 2*PI * (-280 /360) radians = 2*PI * (-70/90) = PI * (-140/90) =PI* (-14/9) = (-14/9) * PI = b (your answer)

OpenStudy (jdoe0001):

@AmyEC so... what's the ANGLE measure of -280 degrees?

OpenStudy (jdoe0001):

well, wait... dohh is already in angle. .ok

OpenStudy (jdoe0001):

\(\bf radians=\cfrac{{\color{brown}{ degrees}}\cdot \pi}{180}\)

OpenStudy (anonymous):

Thank you so much. Do you think you could help with another?

OpenStudy (jdoe0001):

well... easier to post anew, so if we dunno, someone else may, and we can revise each other

OpenStudy (anonymous):

Okay, THank you!